OpenStudy (anonymous):

Convert 45° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

is it pie/6 ?

OpenStudy (ghazi):

1 degree= pi/180 radian...therefore 45 degrees= pi*45/180=pi/4 radians

OpenStudy (anonymous):

oh wow geee thanks !!

OpenStudy (ghazi):

i ain't a geee....you're welcome

OpenStudy (anonymous):

i meant it like ohh jeez thanks

OpenStudy (ghazi):

welcome :)